2016-12-22

MDRS Mission, Sol 3

Here are the science updates, sol records, and commander report from the team at the Hab on the third sol of their mission. Check out Facebook for posts including the photos from each day.


Commander Report – Sol 3

Mission Support,

We’ve been on Mars for 3 sols now, and are forming a strong crew dynamic. Everyone is doing their best to ensure a productive and successful mission. Three of us went on an EVA today (Connor, Anselm, and myself) out to the Dinosaur Quarry to explore the area and get comfortable on the rovers while suited up. It’s been a cold and cloudy day, so the ride out and back was frigid. On our way out to the quarry, we set our pack rat (Bernie) free in the middle of the Martian desert. After returning, we all ate lunch together and discussed our plans for baking holiday treats this weekend.

We are generally staying warm and hydrated; although, while the facility issues have been addressed, we are disappointed with the internet access. Many of us were relying on internet access to complete tasks during our mission, and I personally need wi-fi to complete my research project. We cannot receive email or even navigate to a page on a browser about 90% of the time. This has made evening communications with CapComm, Mission Support, and Shannon Rupert very difficult. We also cannot track our package of mission patches or flight suits without the internet. Due to access settings put in place by the previous Crew Commander, we cannot utilize the other two routers at all. We’ve been monitoring our bandwidth, have turned off all Bluetooth devices, taken turns using our computers, and reset the router numerous times. Nothing has helped. This issue is causing seriously impediments to our mission goals, and we hope there is a way to resolve it with the help of Mission Support.

A Sol Summary, Engineering Report, HSO Report, Science Report, and photos will follow. There is also an EVA Request for Sol 4.

Thank you for the weather report and your support.

Respectfully submitted,

Alison Gibson
Commander, MDRS Crew 171

img_6641-1024x768


Science Log – Sol 3

Greenhab Report
Today began the transplantation of lettuce plants into the Greenhab! We chose our heartiest crop as the test cultivar for GreenHab conditions.  Temperatures have seemed to normalize but today was completely cloud-covered and therefore hostile conditions may return with the sun.  We are leveraging differences in soil and hydroponic systems in order to compare which growing systems may have advantages in an analog Martian habitat. Conveyor plant stages were setup for Red Oak Lettuce today.  Eight developmental stages will be present by the end of the mission. Six were carefully separated from the Rockwool in order to minimize root damage and planted into equal volumes of wet soil.  The hydroponic setup involved creating a hydroponic solution suitable for the broad range of plant life stages present in the conveyor experiment.  This involved starting with tap water and first adding a nutrient solution to get the electrical conductivity (EC) to an appropriate level.  This brought the EC to about 1.9 mS/cm.  The next step was to lower the pH to a slightly basic value.  After calibrating the sensor, the pH was reading 6.06 which was well within the acceptable range.  If all goes well, Green Oak Lettuce and other cultivars will be transplanted tomorrow.

Mars Self Sleep Report Study/Crew Well Being
Last night I tried to get up at 2 am but was not able to do work for long before going back to sleep. I know our plan was to only sleep for three hours at night and nap during the day but we found that the first few days of adjusting to this process are the worst. Almost as bad as flying to Europe or Asia through may time zones and adjusting. We want to raise awareness of the time differences for a Mars mission compared to living on Earth. How will the greater than 24 hr day on Mars translate to astronauts work and sleep patterns? Obviously on Mars, astronauts will be working and fixing things every minute while they are awake. We want to see how we can maximize astronauts’ productivity. As a crew we are going to try and go to sleep earlier one night and start our day by 6 am instead of 8 am like we usually do to see how this improves or reduces productivity. The crew seems to be in great spirits doing their work and surviving here on Mars. Cannot wait to see what SOL 4-13 bring!!

Geology
Today on our EVA we observed many interesting rock outcroppings and geologic formations. It is easy to see the stratified rock layers in exposed hillsides and cliffs. I moved the main weather station here at the HAB from the roof to ground level outside the HAB. The reason for this was that on the roof some of the heat from inside was affecting the temperature measurements. On the next EVA that I go on I plan to put another camera in a location of geologic interest (to be determined) as well as collect the video data from the time-lapse camera that has been sitting near the Hab for a couple days now. This camera will have daytime geology data as well as star exposure. Tomorrow it may snow here so I may give a camera to tomorrow’s EVA team to set up outside somewhere.

I plan to move the weather station inside the GreenHAB within the next couple days so that I can gather accurate weather data first. We want to eventually move it into the GreenHAB in order to use the solar sensor to measure solar flux in the Green HAB compared to outside.

The last time-lapse camera still sits inside the GreenHAB to monitor progress there.

Max/Min
Outdoor Temp – 18 F – 42 F
GreenHab Temp – 47 F – 73 F
Barometer – 29.81 – 29.87 inHg
Wind – 3.5 mph, gust – N/A mph
Solar Rad. Max – 201.7 W/m^2
UV Index – 576 uW/cm^2
Outdoor Humidity – 17% – 41%


Daily Summary Report – Sol 3

SOL: 03
Person filling out Report: Anselm Wiercioch, XO
Summary Title: Longer EVA, beginning greenhab setup
Mission Status: Active. Full crew functional.
Sol Activity Summary: EVA’d to Dinosaur Quarry, set up first conveyor stage in greenhab, planted first 12 plants.
Look Ahead Plan: EVA tomorrow to follow up on initial Quarry recon
Anomalies in work: back side of helmet cracked slightly on rock. Repaired with superglue and zip ties. Not majorly concerning, but helmet henceforth moved to reserve use only.
Weather: High 42F, Low 18F,  wind avg 3.5mph, gust N/A, humidity 17-41%, grey cloudy skies
Crew Physical Status: Slightly stir crazy, but ok.
EVA: Anselm, Geoff, and Alison to Dinosaur Quarry
Reports to be filed:
– Sol Summary
– Commander’s Report
– Science Reports
– 6-8 Photos
– EVA Plan
– Operations Report
Support Requested: Wifi still intermittent.

 

0 Comments

Leave A Comment

Leave a Reply